In a town whose population is 2000, a disease creates an epidemic. The number of people (N) infected (t) days after the disease has begun is given by the function N(t)=2000/(1+19e^-0.6t) a. how many are initially infected with the disease (t=0)? b. how many infected after 2,5,8,12, and 16 days?
Answer by Nate(3500)

N(t)=2000/(1+19e^-0.6t)
N(0)=2000/(1+19(1))
N(0)=2000/(20)
N(0)=100
Initially, 100 people were infected.
--------------------------------------------
N(t)=2000/(1+19e^-0.6t)
N(2)=2000/(1+19e^(-1.2))
I do not have a calculator on me, so I will let you solve the rest of the equation. Also, all you have to do is plug in the numbers into your calculator, and you will have it all solved. I'll let you finish.